y = x³+ 3 x + x³x


1
Expert's answer
2021-02-24T06:07:16-0500

"y=x^3+3x+x^3x"

"\implies y=x^4+x^3+3x"

Slope at any point "(y_1,x_1)" is given by,

"(\frac{dy}{dx})_{(x_1,y_1)}=\frac{d(x^4+x^3+3x)}{dx}_{(x_1,y_1)}"

"=(4x^3+3x^2+3)_{x_1,y_1}"

"=4x_1^3+3x_1^2+3"